Wellcome Technologies PLC, a biotech company, has asked you to evaluate the following project for the production of a new product The firm has spent £200,000 per year on developing this new product for the last 3 years. You are going to charge a fee of £20,000 to undertake the project evaluation. It will be necessary to invest £2.5m in a piece of new machinery immediately. It has been estimated that expected sales are 50000 items of the new product in years 1, 2, 3 and 4 but nothing thereafter. In year 1 you expect items to sell at £50 per item in nominal terms. The nominal total costs of production will be £3,500,000 in year 1. In addition, head office costs will rise because it will be necessary to recruit three new employees on a four-year contract to oversee the project at a starting cost to the firm of £110,000. The project will also be allocated additional costs from the product marketing of £100,000 of head office costs in the first year Question 1 continued/.... Page 1 of 4 Question 1 continued You expect sales revenues and all costs and allocations to rise in line with inflation for years 2, 3 and 4 . Inflation is expected to be 4% per annum. At the end of the fourth year the machinery can be resold for £1,500,000. Working capital is held by the company which is 12.5% of the annual sales. (a) Identify the relevant incremental pre-tax cash flows associated with the project. (30 marks) (b) Assume that Wellcome Technologies PLC is a consistently profitable company. It pays corporation tax at a rate of 25%. For calculating capital allowances for taxation purposes, assume that the new piece of machinery can be depreciated at 20% on a straight-line basis with the first allowance coming immediately (year 0). Identify the relevant post-tax cash flows associated with this project for each year (25 marks) (c) You need to estimate the discount rate for this project. of Wellcome Technologies PLC is 10% and itcan borrow at a pre-tax rate of 3%. It has a target debt to equity ratio (DVE) of 0.6 and is planning on funding this project via a new debt issue. The expected equity return What is the appropriate weighted average cost of capital? What is the NPV of this project? Should Wellcome Technologies PLC go ahead with the project? (25 marks) (d) Why it is important to conduct sensitivity analysis when undertaking capital budgeting? Whch variable do you think the NPV in this case will be most sensitive to?
Answer #1
RELEVANT PRETAX CASH FLOWS:
The following costs are sunk costs needed whether the project is underatken or not
Development costs already incurred £          600,000 (200000*3)
Consultants fees for evaluation of the project £            20,000
The above costs are NOT Relevant Cash lows £          600,000
(a) RELAVANT INCREMENTAL PRE TAX CASH FLOWS:
Sales Revenue in year 1 £      2,500,000 (50000*50)
Sales Revenue in year 2 £      2,600,000 (2500000*1.04)
Sales Revenue in year 3 £      2,704,000 (2600000*1.04)
Sales Revenue in year 4 £      2,812,160 (2704000*1.04)
Cost of production in year 1 £      3,500,000
Cost of production in year 2 £      3,640,000 (3500000*1.04)
Cost of production in year 3 £      3,785,600 (3640000*1.04)
Cost of production in year 4 £      3,937,024 (3785600*1.04)
Year 0 1 2 3 4
A Initial investment on machinery -£     2,500,000
B Sales Revenue £        2,500,000 £      2,600,000 £      2,704,000 £      2,812,160
C Cost of Production -£        3,500,000 -£     3,640,000 -£     3,785,600 -£     3,937,024
D Cost of three new employees -£           110,000 -£         114,400 -£         118,976 -£         123,735
E Additional costs for product marketing -£           100,000 -£         104,000 -£         108,160 -£         112,486
F=B+C+D+E Annual Profit/(Loss) -£        1,210,000 -£     1,258,400 -£     1,308,736 -£     1,361,085
G Working Capital(12.5% of next years sales) £          312,500 £            325,000 £          338,000 £          351,520
H Cash flow due to change in working capital -£         312,500 -£              12,500 -£           13,000 -£           13,520 £          351,520
POST TAX CASH FLOWS:
N Year 0 1 2 3 4
I Dpreciation (20%of 2500000) £          500,000 £            500,000 £          500,000 £          500,000 £          500,000
J=F-I Net Loss including depreciation -£         500,000 -£        1,710,000 -£     1,758,400 -£     1,808,736 -£     1,861,085
K=-0.25*J Tax Savings (25%) £          125,000 £            427,500 £          439,600 £          452,184 £          465,271
L Add :Depreciation (non cas expense) £          500,000 £            500,000 £          500,000 £          500,000 £          500,000
M=J+K+L Annual post tax operating cash flow £          125,000 -£           782,500 -£         818,800 -£         856,552 -£         895,814
P Initial Cash Flow on investment -£     2,500,000
Q Working CapitalCash flow -£         312,500 -£              12,500 -£           13,000 -£           13,520 £          351,520
R=M+P+Q NET POST TAX CASH FLOW -£     2,687,500 -£           795,000 -£         831,800 -£         870,072 -£         544,294
DICOUNT RATE
Ce Cost of Equity 10%
Cd Post tax cost of debt(3*(1-0.25) 2.25%
D=0.6*E
We Weight of Equity Capital =E/(D+E)=1/1.6= 0.63
Wd Weight of Debt Capital =(1-0.63)=                     0.38
Ce*We+Cd*Wd Cost of CAPITAL 7.09%
SUM
PV=R/(1.0709^N) Present Value of NET POST TAX CASH FLOW -£     2,687,500 -£           742,366 -£         725,306 -£         708,449 -£         413,845 -£   5,277,466
NET PRESENT VALUE (NPV) -£     5,277,466
NPV is negative
They should not invest in the project
It is important to conduct sensitivity analysis
to assess the impact of changes in assumptions
of initial investment, Sales and costs